OpenStudy (anonymous):

Find the distance between the points (-7, 5) and (4, 1).

OpenStudy (anonymous):

root {(x2-x1) + (y2-y1)}

OpenStudy (anonymous):

\[\sqrt{\left(x_1-x_2\right){}^2+\left(y_1-y_2\right){}^2}\]

OpenStudy (anonymous):

if u calculate the above, you will find the distance between the two points

OpenStudy (anonymous):

sorry, i missed putting the squares, imran's formula is correct.....